You are going to deposit $26,000 today. You will earn an annual rate of 6.1 percent for 11 years, and then earn an annual rate of 5.5 percent for 14 years. How much will you have in your account in 25 years? O $99.148.20 O $114.250.46 O $105,530.26 O $49.870.37 O $89.871.01
You are going to deposit $2,700 in an account that pays .33 percent interest compounded monthly. How much will you have in 7 years? $3,572.59 O $3,564.40 $3,560.84 $3,543.46 $3,549.12
One advantage of the corporate form of organization is the: ease of formation compared to other organizational forms. unlimited liability for its shareholders. taxation of the corporate profits. double taxation of profits. ability to raise larger sums of equity capital than other organizational forms.
Answer #1

Ans 1) $ 105530.26

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV = PV (1 + r )n
FV = 26000*(1+6.1%)^11 * (1 + 5.5%)^14
FV = 105530.26

Ans 2) $ 3560.84

FV = PV (1 + r )n
FV = 2700*(1+0.33%)^84
FV = 3560.84

Ans 3) Ability to raise large sums or equity capital than other organizational forms.